OpenStudy (anonymous):

Bala had three sons. He had some chocolates which he distributed among them. To his eldest son, he gave 3 chocolates more than half the number of the chocolates with him. To his second eldest son he gives 4 chocolates more than one-third of the remaining number of chocolates with them. To his youngest son he gave 4 chocolates more than one-fourth of the remaining number of chocolates with him. He was left with 11 chocolates. How many chocolates did he initially have?

OpenStudy (anonymous):

(11+4)*4/3= 20 (20+4) *3/2 = 36 (36+3)*2/1 =78 is it 78

OpenStudy (anonymous):

thanks matricked, the answer is correct. but what does the fractions represent?

OpenStudy (anonymous):

move from the last statement to the first (11+4)/(1-1/4) = 20 (20+4) / (1- 1/3) = 36 (36+3)/(1-1/2)=78 is it 78

OpenStudy (anonymous):

thank you, that makes much more sense :)

OpenStudy (anonymous):

yw